# Your Reasoning Benchmark May Not Test Reasoning: Revealing Perception Bottleneck in Abstract Reasoning Benchmarks

### v1

@@ 2 Related Work
Recent progress on ARC/ARC-AGI can be broadly grouped into (i) general-purpose foundation models, including LLMs that serialize grids into text/code and rely on prompting or lightweight adaptation, and VLMs that operate more directly on the visual grid representation; and (ii) tailored ARC solvers that introduce task-specific search, program synthesis, or specialized architectures. The ARC Prize benchmark and its leaderboard further distinguish “base” single-shot model behavior from systems that scale test-time computation, highlighting the central role of inference-time methods in ARC performance (ARC Prize, Inc., 2025). Among tailored approaches, test-time training (TTT) has emerged as a strong general mechanism for few-shot adaptation on ARC, substantially improving over fine-tuned baselines and often combining well with other solver components (Li et al., 2024; Akyürek et al., 2025). In parallel, specialized architectures such as the Hierarchical Reasoning Model (HRM) demonstrate that small recurrent reasoning models can be competitive on ARC (Wang et al., 2025), and follow-up “tiny” recursive variants further push this direction (Jolicoeur-Martineau, 2025). Finally, a very recent concurrent work argues that “ARC is a vision problem” (Hu et al., 2025), reframing ARC as image-to-image translation with a ViT-style backbone and test-time adaptation, which gives an example of a vision-centric route to ARC performance. In this paper, we focus specifically on general-purpose VLMs, aiming to understand how much of their ARC performance is limited by perception rather than reasoning. The recent trend that tailored solutions tend to focus more on enhancing the perception module aligns with our finding. However, our main takeaway is that we should rethink the interpretation of the ARC performance as a measure of general intelligence, as opposed to that we can improve ARC performance by enhancing perception.
@@
The research community has suggested several limitations that complicate treating ARC scores as a clean measure of general reasoning. First, strong results can be driven by solver engineering and compute, such as program-search, heavy verification, and test-time adaptation. Therefore, raw accuracy may reflect an increase in search budget instead of better reasoning. Second, the small hidden evaluation regime and repeated leaderboard-driven iteration can increase the risk of implicit overfitting to quirks of the benchmark (Chollet et al., 2024). Third, ARC tasks are inherently visual, yet many modern pipelines are language-centric; a concurrent work argues that this mismatch can confound conclusions about “reasoning” by bottlenecking on representation, and demonstrates large gains by reframing ARC as image-to-image translation on a continuous canvas” and applying standard vision architectures with test-time training (Hu et al., 2025). In contrast to these benchmark and solver-centric discussions, our study is diagnostic for general-purpose VLMs on ARC-style reasoning: across Mini-ARC, ACRE, and Bongard-LOGO, we hold the underlying tasks fixed and show that explicitly separating perception from reasoning substantially improves performance, showing that performance is largely limited by perceptual fidelity rather than reasoning strength.
@@ 4.1 Datasets
Mini-ARC (Kim et al., 2022) is a reduced-scale variant of ARC with fewer colors, smaller grids, and simplified transformations. It maintains the same input–output format as ARC but with lower perceptual complexity, making it useful for isolating reasoning performance under easier perception.
@@
ACRE (Zhang et al., 2021) is a symbolic visual reasoning benchmark designed for causal induction. Each task is formulated as a classification problem conditioned on demonstrations. Specifically, a task consists of six demonstrations followed by four binary prediction queries. The demonstrations illustrate the presence or absence of certain objects and the corresponding state of a pink board, which may blink (activated), remain unlit (deactivated), or is still underdetermined. Each query presents a new configuration of objects, and the objective is to predict whether the status of the pink board will be activated, deactivated or underdetermined.
@@ 4.3 Setting 2: Stronger-Model Perception
We further conduct experiments corresponding to the second evaluation setting (Stronger-Model Perception) described in Section 3.3, where we use a stronger VLM for the perception stage in the two-stage pipeline. We conduct these experiments on the ACRE dataset, where we replace the LLaVA-1.5 model used in the perception stage with GPT-4o. We also compare against standard one-stage pipelines that use LLaVA-1.5 (weak model) or GPT-4o (strong model). The concrete configurations are summarized in Table 3.
Table 3: Experimental setups in Setting 2 (Stronger-Model Perception) on ACRE. We name the configurations consistent with Table 1 to allow cross reference. The configurations (a) and (b) here are exactly the same as 3(a) and 3(b) in Table 1.
Config Stage Model(s)
(a) S LLaVA-1.5
(b) P+R LLaVA-1.5 (P) + LLaVA-1.5 (R)
(c) P+R GPT-4o (P) + LLaVA-1.5 (R)
(d) S GPT-4o
Results.
As can be seen in Table 4, strengthening the perception module from LLaVA-1.5 to GPT-4o ((b) $\rightarrow$ (c)) yields a substantial improvement, indicating that perception is indeed the dominant bottleneck in this setting. The performance in (c) and that in (d), the standard one-stage pipeline with the strong model GPT-4o, are close, suggesting that perception constitutes the main bottleneck explaining the performance difference between the strong model (GPT-4o) and the weak model (LLaVA-1.5).
Table 4: Success rates (%) on ACRE under Setting 2 (Stronger-Model Perception).
(a) (b) (c) (d)
22.00 34.50 82.50 93.00
4.4 Error Attribution
In order to further validate the hypothesis and gain insights into the causes of errors, we conduct error attribution on the failure cases for different configurations across datasets, following the protocol introduced in Section 3.4. Specifically, for configurations on Mini-ARC and ACRE, we randomly select 50 tasks, and for those on Bongard-LOGO, we randomly select 100 tasks. Among these, some tasks may have been correctly solved by the model, while for the remaining ones we perform detailed error attribution.
Table 5: Error attribution across datasets and experiment configurations. Regarding the settings, for example, “1(a)” refers to the configuration (a) on Mini-ARC, while ‘3(c)‘ refers to the configuration (c) on ACRE.
@@ 4.4 Error Attribution
Results.
As shown in Table 5, perception errors consistently dominate across all setups, indicating that the visual understanding stage is the primary bottleneck. For example, on Mini-ARC, perception errors account for 86.4% of all errors in the setting 1(a) (standard one-stage) and 59.5% in the setting 1(b) (two-stage). A similar trend is observed on Bongard-LOGO, where perception errors represent 65.8% (2(a)) and 37.0% (2(b)) of total errors, and on ACRE, where perception errors reach 76.3% (3(a)) and 68.8% (3(b)). On ACRE with varying P, where different perception modules are compared, the model with LLaVA1.5 still shows 68.8% perception errors, while GPT-4o eliminates perception errors entirely, with all remaining errors in the reasoning stage.
To gain deeper insight into this attribution, we further analyze how error types evolve across experimental configurations—specifically, how perception-related errors transition into reasoning errors (and vice versa) under different configurations, shown in Figure 4. From this analysis, we observe that a substantial portion of perception errors are either eliminated or transformed into downstream reasoning errors, and that the majority of performance gains arise from resolving perception-related errors rather than improving reasoning itself.

### v2

@@ 2 Related Work
Recent progress on ARC/ARC-AGI can be broadly grouped into (i) general-purpose foundation models, including LLMs that serialize grids into text VLMs that operate directly on the visual grid representation; and (ii) tailored ARC solvers that introduce task-specific search, program synthesis, or specialized architectures. For general-purpose foundation models, GPT-5.2 Pro (High) and Gemini 3 Pro (Deep Think) have established a new ceiling, achieving scores of 54.2% and 45.1% respectively in ARC-AGI-2 primarily due to their extended thinking ability (ARC Prize, Inc., 2025). This shift towards scaling test-time computation has crystallized into a new paradigm named “refinement loops,” where systems iteratively optimize solutions against feedback rather than relying on single-shot inference.
Among tailored approaches, test-time training (TTT) has emerged as a strong general mechanism for few-shot adaptation on ARC, substantially improving over fine-tuned baselines and often combining well with other solver components (Li et al., 2024; Akyürek et al., 2025). Alternative neural approaches include the Hierarchical Reasoning Model (HRM) and its recursive variants, which utilize small recurrent architectures (Wang et al., 2025; Jolicoeur-Martineau, 2025), and masked diffusion models, which refine the grid globally to capture structural constraints (Franzen et al., 2025). Additionally, CompressARC explores test-time learning by minimizing description length on the target puzzle (Liao and Gu, 2025). Finally, a very recent concurrent work argues that “ARC is a vision problem” (Hu et al., 2025), reframing ARC as image-to-image translation with a ViT-style backbone and test-time adaptation, which gives an example of a vision-centric route to ARC performance. While this recent trend suggests that enhancing perception boosts performance, our primary goal is to challenge the interpretation of ARC performance as a proxy for general intelligence. Therefore, in this paper, we focus specifically on general-purpose VLMs, aiming to understand how much of their ARC performance is limited by perception rather than reasoning.
@@
The research community has suggested several limitations that complicate treating ARC scores as a clean measure of general reasoning. First, strong performance can result from increased compute, either through larger training budgets or extensive test-time training or searching (Mitchell, 2024). Therefore, gains in accuracy may reflect better fitting to the ARC task distribution rather than improved reasoning. Second, the small hidden evaluation set increases the risk of implicit overfitting to benchmark-specific patterns (Chollet et al., 2024). Third, ARC tasks are inherently visual, yet many recent solutions are language-centric (Li et al., 2024; Akyürek et al., 2025). A concurrent work shows that reframing ARC as an image-to-image translation problem and applying standard vision architectures can achieve near-human performance on ARC-1 (Hu et al., 2025). This result suggests that high ARC scores can be obtained by improving visual representations, without necessarily requiring more general reasoning mechanisms. Our work builds on these studies and examines how perception and reasoning interact when using VLMs to solving ARC tasks.
@@ 4.1 Datasets
Mini-ARC (Kim et al., 2022) is a reduced-scale variant of ARC with fewer colors, smaller grids, and simplified transformations. It maintains the same format as ARC but with lower perceptual complexity, making it useful for isolating reasoning performance under easier perception.
@@
ACRE (Zhang et al., 2021) is a symbolic visual reasoning benchmark designed for causal induction. Each task is formulated as a classification problem conditioned on demonstrations. Specifically, a task consists of six demonstrations followed by four prediction queries. The demonstrations illustrate the presence or absence of certain objects and the corresponding state of a pink board, which may blink (activated), remains unlit (deactivated), or be underdetermined. Each query presents a new configuration of objects, and the objective is to predict the status of the pink board.
@@ 4.3 Setting 2: Stronger-Model Perception
We further conduct experiments corresponding to the second evaluation setting (Stronger-Model Perception) described in Section 3.3, where we use a stronger VLM for the perception stage in the two-stage pipeline. We conduct these experiments on the three datasets mentioned above, where we replace the weaker model used in the perception stage with a stronger model. We also compare against standard one-stage pipelines that use a weak model or a strong model. For Mini-ARC, we consider two different strong models: o1 and Claude-Sonnet-4.5. The concrete configurations are summarized in Table 3.
Table 3: Experimental setups in Setting 2 (Stronger-Model Perception). Configurations (a) and (b) are identical to those in Table 1. Configurations (c) and (d) introduce stronger models either in the perception stage or in a unified one-stage pipeline.
ID Dataset Config Model(s)
1 Mini-ARC (a) GPT-4o (S)
(b) GPT-4o (P) + GPT-4o (R)
(c1) o1 (P) + GPT-4o (R)
(c2) Claude-Sonnet-4.5 (P) + GPT-4o (R)
(d1) o1 (S)
(d2) Claude-Sonnet-4.5 (S)
2 Bongard-LOGO (a) GPT-4o (S)
(b) GPT-4o (P) + GPT-4o (R)
(c) o1 (P) + GPT-4o (R)
(d) o1 (S)
3 ACRE (a) LLaVA-1.5 (S)
(b) LLaVA-1.5 (P) + LLaVA-1.5 (R)
(c) GPT-4o (P) + LLaVA-1.5 (R)
(d) GPT-4o (S)
Table 4: Success rates (%) under Setting 2 (Stronger-Model Perception). Configurations (a)–(d) correspond to those defined in Table 3.
Dataset (a) (b) (c*) (d*)
Mini-ARC 8.05 20.13 (c1) 31.54 (d1) 52.03
(c2) 32.89 (d2) 34.22
Bongard-LOGO 62.00 73.00 80.00 78.00
ACRE 22.00 34.50 82.50 93.00
Table 5: Error attribution across datasets and experiment configurations. Regarding the settings, for example, “1(a)” refers to the configuration (a) on Mini-ARC, while “3(c)” refers to the configuration (c) on ACRE.
@@ 2 Bongard-LOGO (a) GPT-4o (S)
Results.
As can be seen in Table 4, strengthening the perception module ((b) $\rightarrow$ (c*) [^3] We use (c*) as a general reference to (c), (c1), and (c2). Same for (d*) below.) yields consistent and substantial improvement, indicating that perception is indeed the dominant bottleneck in this setting. Furthermore, the performance in (c*) and that in (d*), the standard one-stage pipeline with the strong model, are mostly close [^4] One exception is Mini-ARC with o1, where the gap between (c1) and (d1) remains large. A plausible explanation is that Mini-ARC may have been included in o1’s training data. This possibility is supported by the observation that o1’s performance on Mini-ARC (d1) is markedly higher than that of Claude-Sonnet-4.5 (d2), while its performance on ARC-AGI is inferior to Claude-Sonnet-4.5., suggesting that perception constitutes the main bottleneck explaining the performance difference between the strong model (d*) and the weak model (a).
4.4 Error Attribution
In order to further validate the hypothesis and gain insights into the causes of errors, we conduct error attribution on the failure cases for different configurations [^5] We omitted configurations with o1 or Claude-Sonnet-4.5 as we did not have full reasoning traces in our results. across datasets, following the protocol introduced in Section 3.4. Specifically, for configurations on Mini-ARC and ACRE, we randomly select 50 tasks, and for those on Bongard-LOGO, we randomly select 100 tasks. Among these, some tasks may have been correctly solved by the model, while for the remaining ones we perform detailed error attribution.
Results.
As shown in Table 5, perception errors consistently dominate across all setups, indicating that the visual understanding stage is the primary bottleneck. On Mini-ARC, perception errors account for 86.4% of all errors in setting 1(a) (standard one-stage) and 59.5% in setting 1(b) (two-stage). A similar trend is observed on Bongard-LOGO, where perception errors represent 65.8% (2(a)) and 37.0% (2(b)) of total errors, and on ACRE, where perception errors reach 76.3% (3(a)) and 68.8% (3(b)). On ACRE with varying P, where different perception modules are compared, the model with LLaVA1.5 still shows 68.8% perception errors, while GPT-4o eliminates perception errors, with all remaining errors in reasoning stages.
To gain deeper insight into this attribution, we further analyze how error types evolve across experimental configurations, specifically, how perception-related errors transition into reasoning errors (and vice versa) under different configurations, shown in Figure 4. From the analysis, we observe that a substantial portion of perception errors are either eliminated or transformed into downstream reasoning errors, and that the majority of performance gains arise from resolving perception-related errors rather than improving reasoning.
@@ 5 Conclusion
Limitations
Our study has several limitations. First, the natural-language descriptions used in the perception stage should not be interpreted as an optimal or canonical intermediate representation for abstract reasoning. We adopt language as a convenient and interpretable way to inject generic human perceptual priors and to demonstrate the existence and impact of perceptual bottlenecks in ARC-style benchmarks. We do not claim that natural language is the best representation for isolating reasoning, nor that it preserves all aspects of the original task structure without distortion.
Second, our fine-grained error attribution relies on manual inspection of model outputs and reasoning traces. While this attribution follows a clearly defined, stage-wise protocol and yields consistent patterns across datasets and experimental settings, it involves some subjectivity. We leave the development of scalable and fully automated attribution methods to future work.
Finally, our conclusions are limited to ARC-style, knowledge-light visual abstraction benchmarks and to the vision-language models evaluated in this study. We do not claim that perception is the dominant bottleneck for all multimodal or textual reasoning tasks. Nevertheless, we hope that the diagnostic perspective and methodology introduced in this work can inspire future research on disentangling perception and reasoning in a broader range of reasoning benchmarks and evaluation settings.